# Updating GUT-Scale Pole Higgs Inflation After ACT

### v1

@@ A Generation of the $\mu$ Term of MSSM
where $K_{\rm SUSY}$ can be obtained by expanding $\widetilde{K}_{\ell\rm st}$ in Eq. (5) for low field values with result
$K_{\rm SUSY}=Np\left(|\Phi|^{2}+|\bar{\Phi}|^{2}-\bar{\Phi}\Phi-\bar{\Phi}^{*}\Phi^{*}\right)+|S|^{2}.$ (31)
Also ellipsis in Eq. (30) includes terms which vanish at the SUSY vacuum of Eq. (4) due to the zero v.e.vs of ${H_{u}},{H_{d}}$ and $\widetilde{N}^{c}_{i}$. On the other hand, the contribution into $V_{\rm tot}$ from the soft SUSY-breaking terms read
@@
Given the $\lambda-A_{\rm s}$ condition implied by the rightmost expression in Eq. (20), the derived $\langle{S}\rangle$ turns out to be several orders of magnitude below $m_{3/2}$ since $\lambda\sim 10^{-6}$. On the other hand, the generated $\mu$ parameter from the first term in Eq. (25a) is
$\mu=\lambda_{\mu}\langle{S}\rangle=2\lambda_{\mu}Np\ {\rm a}_{\mu}m_{3/2}/{\lambda}.$ (34)
It can be comparable to $m_{3/2}$ for $N\sim 1$, if $\lambda_{\mu}\sim 10^{-6}$. This prerequisite is welcome in view of Eq. (28) since it assures the stability of the path in Eq. (26) during ppHI.
@@ B Non-Thermal Leptogenesis and $\widetilde{G}$ Constraint
More explicitly, soon after the end of ppHI, the (canonically normalized) inflaton
@@
acquires mass – reducing with increasing $N$ and $p$ – given by
@@
Meanwhile, $\widehat{\delta\phi}$ settles into a phase of damped oscillations abound the minimum in Eq. (4) reheating the universe at a temperature reh ; reh1
$T_{\rm rh}=\left({90\over\pi^{2}g_{\rm rh*}}\right)^{1/4}\left({2(w_{\rm rh}-1)\over 5-3w_{\rm rh}}\widehat{\Gamma}_{\rm\delta\phi}m_{\rm P}\right)^{1/2}$ (37)
with $w_{\rm rh}\simeq 1/3$, $g_{\rm rh*}=228.75$ the MSSM effective number of relativistic degrees of freedom and $\widehat{\Gamma}_{\rm\delta\phi}$ the total decay width of $\widehat{\delta\phi}$
@@
to be consistent with the perturbative bound. The lagrangian terms which describe the decay channels above are similar to those given e.g. in Ref. nmBL .
For $M_{[N^{c}}i]\gg T_{\rm rh}$, the out-of-equilibrium decay of $N^{c}_{i}$ via the second term in Eq. (25a) generates a lepton-number-asymmetry yield which is partially converted through sphaleron effects into a yield of the observed baryon-asymmetry of the universe lept ; zhang
$Y_{B}=-0.35\cdot{5-3w_{\rm rh}\over 4(1-w_{\rm rh})}{T_{\rm rh}\over\widehat{m}_{\rm\delta\phi}}{\widehat{\Gamma}_{\delta\phi\rightarrow N_{i}^{c}}\over\widehat{\Gamma}_{\rm\delta\phi}}\varepsilon_{L}.$ (41)
The simplified formula above can be derived assuming normal hierarchy for $M_{[N^{c}}i]$ – i.e., $M_{[N^{c}}1]\ll M_{[N^{c}}2],M_{[N^{c}}3]$ – and that $\widehat{\delta\phi}$ decays via $\widehat{\Gamma}_{\delta\phi\rightarrow N_{i}^{c}}$ in Eq. (39a) exclusively into $N^{c}_{1}$. Under these circumstances, we can obtain a maximal value for the lepton-number-asymmetry $\varepsilon_{L}$ which is sasa ; dreeslept
@@
$Y_{3/2}\simeq 1.9\cdot 10^{-13}~T_{\rm rh}/{\mbox{\rm EeV}},$ (45)
@@
Fig. 3: Contours yielding the central $Y_{B}$ in Eq. (44) consistently with Eqs. (1), (20), (28), (43) and (46) in the $N-M_{1N^{c}}$ and $N-T_{\rm rh}$ planes for $p=2$, $N_{\rm st}=1/3$ and various $\mu/m_{3/2}$ values indicated on the curves – recall that $1~{\mbox{\rm EeV}}=10^{9}~{\mbox{\rm GeV}}$. The gray [light gray] segments of the lines are allowed by Eq. (46) for $m_{3/2}=10.6~{\mbox{\rm TeV}}$ $m_{3/2}=13.6~{\mbox{\rm TeV}}$].
@@ C Results
The compatibility of the restrictions in Eqs. (28), (43), (44) and (46) together with those from Eqs. (1) and (20) is illustrated in Fig. 3 where we present allowed curves in $N-M_{1N^{c}}$ and $N-T_{\rm rh}$ planes for $p=2$, $N_{\rm st}=1/3$ and $\mu/m_{3/2}=0.01,0.1,1$ and $3$. The chosen $p$ value is very close to the $p$ values which yields the central $n_{\rm s}$ in Eq. (1), as shown in Fig. 2, and the one for $N_{\rm st}$ assists us to overcome the constraint of Eq. (28) without extensive tuning. Moreover, the considered $\mu/m_{3/2}$ ratios are encountered in various recent studies klp ; gabit ; gmssm of the parameter space of MSSM with $m_{3/2}$ values comparable to those considered in Eq. (46). Despite the obtained $\mu$ value is quite low for $\mu/m_{3/2}=0.01$, it is in general still allowed by the LEP lep exclusion limit on the chargino mass.
The ratio $\mu/m_{3/2}$ controls the $\lambda_{\mu}$ value derived by Eq. (34) and via it the contribution of $\widehat{\Gamma}_{\delta\phi\rightarrow H}$ in Eq. (39b) into the $\widehat{\Gamma}_{\rm\delta\phi}$ – see Eq. (38). As expected, the required by Eq. (44) $M_{1N^{c}}$ values increase with $\mu/m_{3/2}$ in order to reinforce the contribution of $\widehat{\Gamma}_{\delta\phi\rightarrow N_{i}^{c}}$ – see Eq. (39a) – into $\widehat{\Gamma}_{\rm\delta\phi}$. The increase of $\mu/m_{3/2}$ is constrained directly by Eq. (28) and indirectly by Eq. (46) since it causes an increase of $T_{\rm rh}$ – see Eq. (37). The gray [light gray] segments of the lines displayed in Fig. 3 are allowed by Eq. (46) for $m_{3/2}=10.6~{\mbox{\rm TeV}}$ $m_{3/2}=13.6~{\mbox{\rm TeV}}$]. As shown from the lower plot in Fig. 3 the lines corresponding to $\mu/m_{3/2}=0.01$ and $0.1$ terminate for low $N$ values due to the upper bound on $T_{\rm rh}$ in Eq. (46). On the other hand, for $\mu/m_{3/2}=1$ and $3$ this termination occurs due to the saturation of the bound on $\lambda_{\mu}$ in Eq. (28). The $\mu/m_{3/2}=3$ line is cut for $N=7.5$ due to the leftmost inequality in Eq. (43). The competition of these two aforementioned constraints causes a progressive reduction of the length of the allowed contours as $\mu/m_{3/2}$ increases beyond $3$ until they shrink to one point for $\mu/m_{3/2}\simeq 5.5$ corresponding to $N=4$, $M_{1N^{c}}=76~{\mbox{\rm EeV}}$ and $T_{\rm rh}=1.73~{\mbox{\rm EeV}}$. Along the curves of Fig. 3 we obtain
@@ VI Conclusions
Prompted by the P-ACT-LB-BK18 data we updated pole HI (i.e. Higgs inflation) sor – or T-Model HI tmhi – by extending our proposal in Ref. actpole . Namely, we adopted the super- and Kähler potentials $W$ and $K$ in Eqs. (2) and (5) which lead to the inflationary potential in Eq. (15) and the kinetic mixing in Eq. (19). Confining $(p,N)$ in the ranges shown in Fig. 2 we achieved a nice covering of the present data – see Fig. 1. Fixing $n_{\rm s}$ to its central value in Eq. (1) we obtained the clear predictions of Eq. (24) regarding our free parameters. As a consequence, our present framework drastically reduces the allowed $(p,N)$ domain found for a gauge-singlet inflaton in Ref. actpole . It is gratifying that the resulting $r$ values for $(p,N)$’s of order unity could be observationally accessible in the foreseeable future. We also specified a post-inflationary completion, based on the extra $W$ and $K$ terms in Eqs. (25a) and (25b), which offers a nice solution to the $\mu$ problem of MSSM and allows for baryogenesis via nTL with $M_{1N^{c}}$ in the range $(10-76){\mbox{\rm EeV}}$ provided that $\mu/m_{3/2}\leq 5.5$.

### v2

@@ A Generation of the $\mu$ Term of MSSM
where $K_{\rm SUSY}$ can be obtained by expanding $\widetilde{K}_{\ell\rm st}+\Delta K$ – see Eqs. (4) and (24b) – for low field values with result
$K_{\rm SUSY}=Np\left(|\Phi|^{2}+|\bar{\Phi}|^{2}-\bar{\Phi}\Phi-\bar{\Phi}^{*}\Phi^{*}\right)+|S|^{2}+|X^{\alpha}|^{2}.$ (30)
Also ellipsis in Eq. (29) includes terms which vanish at the SUSY vacuum due to the zero v.e.vs of ${H_{u}},{H_{d}}$ and $\widetilde{N}^{c}_{i}$. On the other hand, the contribution into $V_{\rm tot}$ from the soft SUSY-breaking terms read
@@
The generated $\mu$ parameter from the first term in Eq. (24a),
$\mu=\lambda_{\mu}\langle{S}\rangle=2\lambda_{\mu}Np\ {\rm a}_{\mu}m_{3/2}/{\lambda},$ (33)
can be comparable to $m_{3/2}$ for $N\sim 1$, if $\lambda_{\mu}$ is of the same order of magnitude with $\lambda$. Since $\lambda\sim 10^{-6}$, due to $\lambda-A_{\rm s}$ condition implied by the rightmost expression in Eq. (IV) – see Ref. actpole –, the resulting $\lambda_{\mu}$ can become compatible with Eq. (27) assuring, thereby, the stability of the path in Eq. (25) during ppHI.
@@ B Non-Thermal Leptogenesis and $\widetilde{G}$ Constraint
More explicitly, when ppHIis over, the (canonically normalized) inflaton
@@
continues to roll down towards the SUSY vacuum, Eq. (3), where it acquires mass – reducing with increasing $N$ and $p$ – given by
@@
After a brief stage of tachyonic preheating preheating which does not lead to significant particle production garcia , $\widehat{\delta\phi}$ settles into a phase of damped oscillations initially around zero – where $V_{\rm HI}$ has a maximum – and then around one of the minima of $V_{\rm HI}$ similarly to the picture described in Appendix B of Ref. hi5 for the non-minimal HI. Whenever $\widehat{\delta\phi}$ passes through zero, several particles (mostly $A_{BL}$ bosons) may be produced via the mechanism of instant preheating instant . This process becomes more efficient as $N$ decreases since the amplitude of the first oscillation
$\phi_{\rm mx}\simeq\sqrt{6/(16+3Np)}m_{\rm P}$ (36)
increases, increasing thereby the frequency of the violation of the adiabaticity criterion instant ; garcia . Given that the efficiency of this mechanism requires a rather large (let say 507050-70) number of passages of $\widehat{\delta\phi}$ through zero, we intuitively expect that our estimates below are more accurate for $N>0.01$ which ensures a less frequent passage of $\widehat{\delta\phi}$ through zero weakening, thereby, the effects from preheating.
Nonetheless the standard perturbative approach to the $\widehat{\delta\phi}$ decay provides a very efficient total decay rate
@@
to be consistent with the perturbative bound. The lagrangian terms which describe the decay channels above are similar to those given e.g. in Ref. nmBL . Note that three-particle decay of $\widehat{\delta\phi}$ is suppressed as noted there and possible perturbative decay into $A_{BL}$ bosons is kinetically blocked. Thanks to $\widehat{\Gamma}_{\rm\delta\phi}$ the universe is reheated at a temperature reh ; reh1
$T_{\rm rh}=\left({90\over\pi^{2}g_{\rm rh*}}\right)^{1/4}\left({2(w_{\rm rh}+1)\over 5-3w_{\rm rh}}m_{\rm P}\widehat{\Gamma}_{\rm\delta\phi}\right)^{1/2}$ (40)
with $g_{\rm rh*}=228.75$ the MSSM effective number of relativistic degrees of freedom.
For $M_{[N^{c}}i]\gg T_{\rm rh}$, the out-of-equilibrium decay of $N^{c}_{i}$ via the second term in Eq. (24a) generates a $L$-asymmetry yield which is partially converted through sphaleron effects into a yield of the observed $B$ asymmetry of the universe lept ; zhang
$Y_{B}=-0.35\cdot 2\cdot{5-3w_{\rm rh}\over 4(1-w_{\rm rh})}{T_{\rm rh}\over\widehat{m}_{\rm\delta\phi}}{\widehat{\Gamma}_{\delta\phi\rightarrow N_{i}^{c}}\over\widehat{\Gamma}_{\rm\delta\phi}}\varepsilon_{L}.$ (41)
The simplified formula above can be derived assuming normal hierarchy for $M_{[N^{c}}i]$ – i.e., $M_{[N^{c}}1]\ll M_{[N^{c}}2],M_{[N^{c}}3]$ – and that $\widehat{\delta\phi}$ decays via $\widehat{\Gamma}_{\delta\phi\rightarrow N_{i}^{c}}$ in Eq. (38a) exclusively into $N^{c}_{1}$. Under these circumstances, we can obtain a maximal value for the $L$ asymmetry $\varepsilon_{L}$ which is sasa ; dreeslept
@@
$Y_{3/2}\simeq 2.9\cdot 10^{-13}(T_{\rm rh}/{\mbox{\rm EeV}})/(3-w_{\rm rh}),$ (45)
@@
Fig. 3: Contours yielding the central $Y_{B}$ in Eq. (44) consistently with Eqs. (1), (IV), (27), (43) and (46) in the $N-M_{1N^{c}}$ and $N-T_{\rm rh}$ planes for $p=2$, $N_{\rm st}=1/3$ and various $\mu/m_{3/2}$ values indicated on the curves – recall that $1~{\mbox{\rm EeV}}=10^{9}~{\mbox{\rm GeV}}$. The gray [light gray] segments of the lines are allowed by Eq. (46) for $m_{3/2}=10.6~{\mbox{\rm TeV}}$ $m_{3/2}=13.6~{\mbox{\rm TeV}}$].
@@ C Results
The compatibility of the restrictions in Eqs. (27), (43), (44) and (46) together with those from Eqs. (1) and (IV) is illustrated in Fig. 3 where we present allowed curves in $N-M_{1N^{c}}$ and $N-T_{\rm rh}$ planes for $p=2$, $N_{\rm st}=1/3$ and $\mu/m_{3/2}=0.01,0.1,1$ and $3$. The chosen $p$ value is very close to the $p$ values which yields the central $n_{\rm s}$ in Eq. (1), as shown in Fig. 2, and the one for $N_{\rm st}$ assists us to overcome the constraint of Eq. (27) without extensive tuning. Moreover, the considered $\mu/m_{3/2}$ ratios are encountered in various recent studies klp ; gabit ; gmssm of the parameter space of MSSM with $m_{3/2}$ values comparable to those considered in Eq. (46). Despite the obtained $\mu$ value is quite low for $\mu/m_{3/2}=0.01$, it is in general still allowed by the LEP lep exclusion limit on the chargino mass.
The ratio $\mu/m_{3/2}$ controls the $\lambda_{\mu}$ value derived by Eq. (33) and via it the contribution of $\widehat{\Gamma}_{\delta\phi\rightarrow H}$ in Eq. (38b) into the $\widehat{\Gamma}_{\rm\delta\phi}$ – see Eq. (37). As expected, the required by Eq. (44) $M_{1N^{c}}$ increases with $\mu/m_{3/2}$ in order to reinforce the contribution of $\widehat{\Gamma}_{\delta\phi\rightarrow N_{i}^{c}}$ – see Eq. (38a) – into $\widehat{\Gamma}_{\rm\delta\phi}$. The increase of $\mu/m_{3/2}$ is constrained directly by Eq. (27) and indirectly by Eq. (46) since it causes an increase of $T_{\rm rh}$ – see Eq. (40). The gray [light gray] segments of the lines displayed in Fig. 3 are allowed by Eq. (46) for $m_{3/2}=10.6~{\mbox{\rm TeV}}$ $m_{3/2}=13.6~{\mbox{\rm TeV}}$]. As shown from the lower plot in Fig. 3 the lines corresponding to $\mu/m_{3/2}=0.01$ and $0.1$ terminate for low $N$ values due to the upper bound on $T_{\rm rh}$ in Eq. (46). On the other hand, for $\mu/m_{3/2}=1$ and $3$ this termination occurs due to the saturation of the bound on $\lambda_{\mu}$ in Eq. (27). The $\mu/m_{3/2}=3$ line is cut for $N=7.5$ due to the leftmost inequality in Eq. (43). The competition of these two aforementioned constraints causes a progressive reduction of the length of the allowed contours as $\mu/m_{3/2}$ increases beyond $3$ until they shrink to one point for $\mu/m_{3/2}\simeq 5.5$ corresponding to $N=4$, $M_{1N^{c}}=76~{\mbox{\rm EeV}}$ and $T_{\rm rh}=1.73~{\mbox{\rm EeV}}$. Along the curves of Fig. 3 we obtain
@@ VI Conclusions
Prompted by the P-ACT-LB-BK18 data we updated pole sor – or T-Model tmhi – HI (i.e. Higgs inflation) by extending our proposal in Ref. actpole . Namely, we adopted the super- and Kähler potentials $W$ and $K$ in Eqs. (2) and (4) which lead to the inflationary potential in Eq. (14) and the kinetic mixing in Eq. (18). Moreover we required that the Higgs-inflaton assumes ultimately its v.e.v predicted by the gauge unification within MSSM. This requirement influenced via Eq. (21) the parameter space of ppHI and via Eqs. (35) and (39) the post-inflationary scenario. Confining $(p,N)$ in the ranges shown in Fig. 2 we achieved a nice covering of the present data – see Fig. 1. Fixing $n_{\rm s}$ to its central value in Eq. (1) we obtained the clear predictions of Eq. (23) regarding our free parameters. As a consequence, our present framework drastically reduces the allowed $(p,N)$ domain found for a gauge-singlet inflaton in Ref. actpole . It is gratifying that the resulting $r$ values for $(p,N)$’s of order unity could be observationally accessible in the foreseeable future. We also specified a post-inflationary completion, based on the extra $W$ and $K$ terms in Eqs. (24a) and (24b), which offers a nice solution to the $\mu$ problem of MSSM and allows for baryogenesis via nTL with $M_{1N^{c}}$ in the range $(10-76)~{\mbox{\rm EeV}}$ provided that $\mu/m_{3/2}\leq 5.5$.